OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. C(0, 4), m = 0

OpenStudy (anonymous):

\[y-4=0(x-0)\]

OpenStudy (anonymous):

so what will be the answer? @BellaRain

OpenStudy (anonymous):

y+x=4? @jinxhead20

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

Yay! Thank you!

OpenStudy (anonymous):

no prob